the difference between 2 numers is 1300.Nowing that 2/7 of the biggest number represents 7/18 of the second number, find the 2 numers !!!!!!!!!!
I REALLY NEED HELP !!!!!! PLZ!

Answer by amine(11)

we suppose a and b are these number then
a-b=1300
2/7a=7/18b => a=49/36b we use it in the first (49/36-1)b=1300 =>b=3600 and a=4900